A large container holds 14 litres of a solution that is 25% antifreeze, the remainder being water. How many litres of antifreeze must be added to the container to make a solution that is 30% antifreeze?
Please could you help me I am seriously confused.
25% of 14 = 3.5 remainder is 10.5l of water
Please help?

Answer by scott8148(6628)

14(25%) + x (100%) = (14 + x)(30%)

35 + 10x = 42 + 3x

7x = 7 ___ x = 1

adding one liter of antifreeze gives 4.5 liters in 15 liters of solution or 30%